OpenStudy (anonymous):

Find dy/dx, if y = 10^tan(u) * log base 5 of (sec^2 u^3)

OpenStudy (anonymous):

there's no x! :D

OpenStudy (anonymous):

right, dy/du then

OpenStudy (anonymous):

dy/dx=0

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

hmm i am a little slow. product rule and \[\frac{d}{du}10^{\tan(u)}=10^{\tan(u)}\sec^2(u) \ln(10)\] so far so good?

OpenStudy (anonymous):

@imran what constant is this?

OpenStudy (anonymous):

I am finding dy/dx , not dy/du

OpenStudy (anonymous):

Satellite, you used the product rule?

OpenStudy (anonymous):

@ imran lol @soda i didn't do anything yet except take the derivative of \[10^{\tan(u)}\]

OpenStudy (anonymous):

you still have to take the derivative of \[\log_5(\sec^2(u^3))\] which will require some chain rule business, but before we do that lets rewrite it as \[2\log_5(\sec(u^3))\] to make life a tiny bit easier

OpenStudy (anonymous):

good god this is a pain. your math teacher must hate you

OpenStudy (anonymous):

How did you find the derivative of \[10^\tan u\] to be \[10^\tan u * \sec^2 u \ln 10\]?

OpenStudy (anonymous):

chain rule and the fact that \[\frac{d}{dx}10^x=10^x\ln(10)\]

OpenStudy (anonymous):

or more generally \[\frac{d}{dx}b^x=b^x\ln(b)\]

OpenStudy (anonymous):

so where does the \[\sec ^2 u \] come from?

OpenStudy (anonymous):

the ln(b) is where it comes from, right?

OpenStudy (anonymous):

it is the derivative of \[\tan(u)\] so by the chain rule we get \[\frac{d}{du}10^{\tan(u)}=10^{\tan(u)}\ln(10)\times \frac{d}{du}\tan(u)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now we still have to take the derivative of \[2\log_5(\sec(u^3))\] again a big chain rule problem

OpenStudy (anonymous):

would you change \[2\log_{5}(\sec u^3)\] to exponential form?

OpenStudy (anonymous):

are you sure it is not dy/dx?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the derivative of \[\log_b(x)\] is \[\frac{1}{\ln(b)x}\] so here we will get \[\frac{2}{\ln(5)\sec(u^3)}\times \frac{d}{du}\sec(u^3)\] \[\frac{2\sec(u^3)\tan(u^3)\times 3u^2}{\ln(5)\sec(u^3)}\]

OpenStudy (anonymous):

or after canceling you get \[\frac{6u^2\tan(u^3)}{\ln(5)}\]

OpenStudy (anonymous):

You made that easy to understand Sate, thanks!

OpenStudy (anonymous):

and we are still not done. because you have to use the product rule for this one \[(fg)'=f'g+g'f\] with \[f(u)=10^{\tan(u)}, f'(u)=10^{\tan(u)}\sec^2(u) \ln(10)\] \[g(u)=2\log_5(\sec(u^3)),g'(u)=\frac{6u^2\tan(u^3)}{\ln(5)}\]

OpenStudy (anonymous):

i will let you do that on your own

OpenStudy (anonymous):

Yeah, I can handle the product having the derivatives, I just had trouble trying to find the derivatives.

OpenStudy (anonymous):

good luck and yw